OpenStudy (anonymous):

Solve: 40*x + 52 = 32*x + 48

OpenStudy (anonymous):

i need help ty

OpenStudy (texaschic101):

40x + 52 = 32x + 48 -- subtract 32x 40x - 32x + 52 = 48 -- subtract 52 40x - 32x = 48 - 52 -- combine like terms 8x = - 4 -- divide by 8 x = -4/8 reduces to -1/2

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

nope i have more problems thou

OpenStudy (texaschic101):

I can help